on her last two math tests. leah had scores of 83 and 92. what must leah score on the next test to average a 90 on all three test
Answer by richwmiller(17219)

two ways to do this
90+7-2=95
other longer way
(83+92+x)/3=90
(x+175)/3=90
x+175=270
x=95